OpenStudy (anonymous):

I need help with Factorization 3

OpenStudy (stormfire1):

ask away..

OpenStudy (anonymous):

3 * 1 = 3 :-)

OpenStudy (anonymous):



OpenStudy (anonymous):

No I know simple multiplication I need help with factorization 3 ( you know like algerbra 2 )

OpenStudy (anonymous):

OK... what is your question?

OpenStudy (anonymous):

to factor that type of polynomial i need to do the cross multiple cation method

OpenStudy (anonymous):

2x squared + 5x +3x

OpenStudy (stormfire1):

\[2x^2 + 5x + 3x = 2x^2 + 8x\]

OpenStudy (anonymous):

2x^2 + 5x + 3x = 2x^2 + 8x = 2x(x + 4)

OpenStudy (anonymous):

i thought it was ( x + 1 ) ( 2x +3 ) ?

OpenStudy (anonymous):

that would make 2x^2 + 5x + 3. Did you type your original in wrong?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you said "2x squared + 5x +3x"